OpenStudy (anonymous):

A laptop computer is purchased for $1850 . After each year, the resale value decreases by 35% . What will the resale value be after 4 years? Use the calculator provided and round your answer to the nearest dollar.

OpenStudy (mikwind):

\[V=V_0(1-loss)^a\] , where a is the number of years, and loss is the percentage of lost value per year

OpenStudy (anonymous):

2472221600

OpenStudy (anonymous):

@welshfella can you help me

OpenStudy (mikwind):

Lets simplify this a little: After 1 year the new value is : 1850*(100%-35%)= 1850*0.65=1202.5, after another year it is 1202.5*(0.65)=781,625, after another year it is 781,625*0.65=... Every year you multiply by 0.65

OpenStudy (welshfella):

just carry on where mikwind finished